OpenStudy (anonymous):

Simplify, this just hurts my head 0.o 120m^16/ 45m^8

OpenStudy (anonymous):

\[\frac{ 120m ^{16} }{ 45m^{8} }\] is the same as \[\frac{ 120 }{ 45 }.\frac{m ^{16} }{ m^{8} }\] so first simplify the constant \[\frac{ 120 }{ 45 }=\frac{ 8 }{ 3 }\] then simplify the powers \[\frac{ m ^{16} }{ m^{8} }=m^8\] using \[\frac{ m ^{a} }{ m^{b} }=m ^{a-b} \]thus, \[\frac{ 120m ^{16} }{ 45m^{8} }=\frac{ 8m ^{8} }{ 3 }\]